# A Generative Modeling / Physics-Informed Neural Network Approach to Random Differential Equations

### v1

@@ 2.4 Variational Approximation of the Target Law
where $d_{\mathcal{P}_{2}^{\gamma}(\tilde{\mathcal{H}})}$ and $d_{\mathcal{P}_{2}^{\gamma}(\overline{\mathcal{H}})}$ are appropriate metrics defined over the space of probability measures $\mathcal{P}_{2}^{\gamma}\left(\tilde{\mathcal{H}}\right)$ and $\mathcal{P}_{2}^{\gamma}\left(\overline{\mathcal{H}}\right)$, respectively. The variational problem then becomes:
$\mu_{\theta^{*}}\,\,\in\,\,{\operatorname*{arg\,min}}_{\mu_{\theta}\in\mathcal{V}_{\mathcal{N}}^{*}}\,\,d_{\mathcal{P}_{2}^{\gamma}(\tilde{\mathcal{H}})}\left(A\odot\nu_{u},A\odot\mu_{\theta}\right)+d_{\mathcal{P}_{2}^{\gamma}(\overline{\mathcal{H}})}\left(B\odot\nu_{u},B\odot\mu_{\theta}\right)\,.$ (11)
@@ 3.1 Fully Network Based Neural Measure Spaces
$\mathcal{V}_{\mathcal{N}}:=\left\{g_{\theta}:\mathbb{R}^{d_{in}}\times\mathbb{R}^{d_{ran}}\to\mathbb{R}^{s}:g_{\theta}(x,t;\xi):=\ C_{L}\circ\sigma\circ C_{L-1}\cdots\circ\sigma\circ C_{1}(x,t;\xi)\text{ for }\theta\in\Theta\right\}\,.$
@@
$C_{k}y=W_{k}y+b_{k},\qquad\text{where }W_{k}\in\mathbb{R}^{d_{k+1}\times d_{k}},b_{k}\in\mathbb{R}^{d_{k+1}}\,,$ (12)
@@
$X_{\theta}(\xi)=g_{\theta}\left(\cdot;\xi\right),\qquad\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}\,.$
@@ 3.2 PCE-NN Based Neural Measure Spaces
To this end, consider a Polynomial Chaos Expansion (PCE)-type basis, [PCE_karniadakis], $\left\{\phi_{n}\right\}_{n\in\mathbb{N}}\subset L^{2}(\gamma)$ satisfying:
$\forall q\in L^{2}(\gamma;\mathcal{H}),\quad\exists\,!\,\left\{a_{n}\right\}_{n\in\mathbb{N}}\subset\mathcal{H}:\quad q(x,t;\xi)=\sum_{n\in\mathbb{N}}a_{n}(x,t)\phi_{n}(\xi)\vspace{-0.3cm}\,,$ (13)
@@
$\mathcal{V}_{\mathcal{N}}:=\left\{g_{\theta}:\mathbb{R}^{d_{input}}\to\mathbb{R}^{d_{out}K}:\ g_{\theta}=g_{\theta}(x,t)\text{ for }\theta\in\Theta\right\}\,,$
and the structure of the neural network functions $g_{\theta}(x,t)$ is as before, but now affecting only $x,t$ variables. Then the corresponding form of neural measures is:
$\mathcal{V}_{\mathcal{N}}^{*}=\left\{\left[X_{\theta}\right]_{\!\#}\gamma\in\mathcal{P}_{2}^{\gamma}\left(\,\,\mathcal{H}\,\,\right):\text{ where }X_{\theta}(\xi)=\sum_{n=1}^{K}g_{\theta}^{(n)}\,\phi_{n}(\xi)\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}\right\}\,.$
@@ 3.3 Galerkin-NN Based Neural Measure Spaces
Consider appropriate a given finite dimensional Galerkin subspace of $\mathcal{H}$, denoted by $\mathcal{H}_{h},$ let $M$ be its dimension and $\left\{\psi_{n}\right\}_{1}^{M}$ its basis. Then we consider the elements of $L^{2}(\gamma;\mathcal{H})$
$q(x,t;\xi)=\sum_{n=1}^{M}b_{n}(\xi)\psi_{n}(x,t)\,.$ (14)
@@
$\mathcal{V}_{\mathcal{N}}:=\left\{g_{\theta}:\mathbb{R}^{d_{ran}}\to\mathbb{R}^{M}:g_{\theta}=g_{\theta}(\xi)\ \ \text{ for }\theta\in\Theta\right\}\,,$
and the structure of the neural network functions $g_{\theta}(\xi)$ is as before, but now affecting only the $\xi$ variable. Then the corresponding form of neural measures is:
@@ 4.1 Loss based on Wasserstein distance
Within the framework the present paper it is particularly convenient to consider Wasserstein distances. A key observation is that when one of the measures to be used in the loss is atomic, i.e., $\delta_{g},$ then the set of plans is reduced to $\mu\times\delta_{g},$ i.e., $\Pi(\mu,\delta_{g})=\{\mu\times\delta_{g}\}$, see (5.2.12) in [AGS:2005]. Assuming $g\in\mathcal{H}$ and $\delta_{g}\in\mathcal{P}_{p}(\mathcal{H})$, we then have
$\begin{split}W_{p}(\mu,\delta_{g})^{p}=&\inf_{\pi\in\Pi(\mu,\delta_{g})}\int_{\mathcal{H}}\int_{\mathcal{H}}\left\|x-y\right\|_{\mathcal{H}}^{p}\pi({\rm d}x,{\rm d}y)=\int_{\mathcal{H}}\int_{\mathcal{H}}\left\|x-y\right\|_{\mathcal{H}}^{p}\mu({\rm d}x)\delta_{g}({\rm d}y)\\ =&\int_{\mathcal{H}}\left\|x-g\right\|_{\mathcal{H}}^{p}\mu({\rm d}x)=\mathbb{E}_{X\sim\mu}\left\|X-g\right\|_{\mathcal{H}}^{p}\end{split}\,.$ (15)
We distinguish two cases. In case where $\mathcal{A}\mathcal{U}(\xi)$ and $\mathcal{B}\mathcal{U}(\xi)$ are constant with respect to $\xi$, we can easily check exists $f\in\tilde{\mathcal{H}}$ and $g\in\overline{\mathcal{H}}$ such that
@@
$\mu_{\theta^{*}}\in{\operatorname*{arg\,min}}_{\mu_{\theta}\in\mathcal{V}_{\mathcal{N}}^{*}}\,\,W_{p}(A\odot\mu_{\theta},\delta_{f})+W_{p}(B\odot\mu_{\theta},\delta_{g})\,.$
@@
$\tilde{A}_{\xi}X_{\theta}(\xi):=\mathcal{A}\mathcal{X}_{\theta}(\xi)-\mathcal{A}\mathcal{U}_{\theta}(\xi),\quad\text{ and }\quad\tilde{B}_{\xi}X_{\theta}(\xi):=\mathcal{B}\mathcal{X}_{\theta}(\xi)-\mathcal{B}\mathcal{U}_{\theta}(\xi)\,,$
@@
$d_{2,\gamma}\left(A\odot\nu_{u},\,A\odot\mu_{\theta}\right)^{2}+d_{2,\gamma}\left(B\odot\nu_{u},\,B\odot\mu_{\theta}\right)^{2}=\left\|\mathcal{A}\mathcal{U}-\mathcal{A}\mathcal{X}_{\theta}\right\|_{L^{2}(\gamma;\tilde{\mathcal{H}})}^{2}+\left\|\mathcal{B}\mathcal{U}-\mathcal{B}\mathcal{X}_{\theta}\right\|_{L^{2}(\gamma;\overline{\mathcal{H}})}^{2}\,.$
In fact, recall that both $\nu_{u},\,\mu_{\theta}\in\mathcal{P}_{2}^{\gamma}(\mathcal{H})$. Additionally, by hypothesis (7), it follows that $A\odot\nu_{u},\,A\odot\mu_{\theta}\in\mathcal{P}_{2}^{\gamma}(\tilde{\mathcal{H}})$ and $B\odot\nu_{u},\,B\odot\mu_{\theta}\in\mathcal{P}_{2}^{\gamma}(\overline{\mathcal{H}})$.
@@ 4.2 Other metrics and divergences
It is possible, however, to utilise alternative divergences or metrics to define the loss functional. We briefly review some potential choices below, see, eg., [Katsoul_pantazis2022, renyi1961measures, muller1997integral, sriperumbudur2009integral, birrell2022f, Villani_OT_book_2009, sen2011introduction].
@@ 4.2.1 Divergences and Metrics: Definitions and Relationships
Divergences generalise metrics by dropping symmetry and the triangle inequality, while retaining non-negativity and vanishing only when the inputs are equal. A key family is the f-divergences: given probability measures $\mu\ll\nu$ and a convex function $f:\mathbb{R}_{+}\to\mathbb{R}\cup\left\{+\infty\right\}$ with $f(1)=0$, the flimit-$f-$divergence is defined as
@@
    $d_{\mathcal{F}}(\mu,\nu):=\sup_{f\in\mathcal{F}}\left|\mathbb{E}_{X\sim\mu}f(X)-\mathbb{E}_{Y\sim\nu}f(Y)\right|\,.$ (16)
@@
    Given a probability measure $\lambda\in\mathcal{P}(\mathcal{H})$, we define $\mathcal{P}_{\lambda}(\mathcal{H})\subset\mathcal{P}(\mathcal{H})$ to be the collection of probability measures that are absolutely continuous with respect to $\lambda$, ie. $\mu\ll\lambda$. Let $\mu,\,\nu\in\mathcal{P}_{\lambda}(\mathcal{H})$ with Radon-Nikodym derivatives $f_{\mu},\,f_{\nu}$ with respect to $\lambda$. Then TV distance between $\mu,\,\nu$ can be expressed as:
    $d_{TV}(\mu,\nu):=\frac{1}{2}\left\|f_{\mu}-f_{\nu}\right\|_{L^{1}(\lambda)}.$
@@ 4.2.2 Approximating Distances in Finite-Dimensional Spaces
Given that our spaces, $V_{\Theta}=V_{\mathcal{N}}^{*}$, have a finite number of parameters, we can employ techniques developed in approximating distances of distributions in finite dimensions to compute the actual loss function. Indicative such methods include, GANs and computational optimal transport algorithms, [goodfellow2014generative, Peyre:CompOT:2019].
For approximating Wasserstein distances, one may use Sinkhorn Algorithm which introduces entropic regularisation for efficient computation and other approaches detailed in [Peyre:CompOT:2019]. These approaches exploit the geometric and variational structure of optimal transport theory and are particularly relevant in generative modelling and distributional approximation tasks.
Generative Adversarial Networks (GANs) provide a flexible variational framework for approximating probability distributions by minimising statistical distances, notably Integral Probability Metrics (IPMs) and, under certain conditions, flimit-$f-$divergences.
@@
In the special case where the class of test functions $\mathcal{F}$ associated with a particular IPM (eg. as in equation (16)) can be approximated by the class of discriminators $\overline{\mathcal{V}}_{\mathcal{N}}$. Then, this min-max formulation can be viewed as minimising IPM between $\mathbb{P}_{r}$ and generators distribution. Specifically, we define the class of neural measures:
$\mathcal{V}_{\mathcal{N}}^{*}:=\left\{\left[g_{\theta}\right]_{\!\#}\pi_{0}:\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}\right\}\,.$
@@
$\operatorname*{arg\,min}_{\mu_{\theta}\in\mathcal{V}_{\mathcal{N}}^{*}}d_{\mathcal{F}}(\mathbb{P}_{r},\mu_{\theta}),\quad\text{ where }d_{\mathcal{F}}\text{ is the IPM induced by the class }\mathcal{F}.$
@@ 5 Numerical results
following [Ziyin:2020]. This choice consistently yielded better convergence and expressiveness than the traditional hyperbolic tangent activation, especially for problems involving both periodic and non-periodic solutions. The parameter $a$ is specific to each layer and is optimized jointly with the network weights during training.
@@
All experiments were conducted using a unified codebase, which is publicly available at [arampatzis:2024], ensuring reproducibility and ease of extension to other problem settings.
@@ 5.1 Bistable ordinary differential equation
$\frac{\mathrm{d}u}{\mathrm{d}t}=-r(u-1)(2-u)(y-3)\,,\quad t\in[0,8]\,,$ (17)
@@
$u(0)=u_{0}\,.$ (18)
The paramaters $u_{0},r$ follow uniform distributions, $y_{0}\sim\mathcal{U}(0,4)$ and $r\sim\mathcal{U}(0.8,1.2)$. The equation has two stable ($u=1$, $u=3$) and one unstable ($u=2$) equilibrium solutions.
Figure 1: Training and testing loss over time during the training of the PINN architecture for the bistable ODE (17).
(a) Histogram comparison.
(b) Wasserstein distance over time.
Figure 2: Histograms at time $t=2$ and Wasserstein distance between the histograms over the time interval $[0,8]$ of the solution of the bistable ODE (17) using an implicit numerical solver and the PINN architecture for different parameter values.
In Figure 1, we plot the training and testing loss functions throughout the optimization. The spatial-temporal domain is resampled every 50 iterations, as are the parameters, which are drawn from their predefined uniform distribution. Each resampling uses 100 random samples from the domain. After each resampling step, the loss typically spikes due to the shift in training data but continues to decrease overall as training progresses. The testing loss begins to rise after around 400 iterations, indicating that the network starts to overfit the training data.
In Figure 2, we plot histograms of the solution to the bistable ODE (17). The reference solution is obtained using an implicit numerical solver. We observe good qualitative agreement between the reference and the PINN-based histograms. However, the Wasserstein distance between the two solutions increases over time. This is expected, as the solution’s distribution gradually concentrates into two Dirac delta peaks at $u=1$ and $u=3$. This sharp concentration makes convergence challenging, since the Wasserstein distance is highly sensitive to samples that have not fully collapsed onto the Dirac masses.

### v2

@@ 2.4 Variational Approximation of the Target Law
[…]
(11) $\mu_{\theta^{*}}\in\mathop{\mathrm{arg\,min}}_{\mu_{\theta}\in\mathcal{V}_{\mathcal{N}}^{*}}d_{\mathcal{P}_{2}^{\gamma}(\tilde{\mathcal{H}})}(A\odot\nu_{u},A\odot\mu_{\theta})+d_{\mathcal{P}_{2}^{\gamma}(\overline{\mathcal{H}})}(B\odot\nu_{u},B\odot\mu_{\theta}).$
@@ 3.1 Fully Network Based Neural Measure Spaces
$\mathcal{V}_{\mathcal{N}}\coloneqq\left\{\begin{aligned} g_{\theta}:\mathbb{R}^{d_{\text{in}}}\times\mathbb{R}^{d_{\text{ran}}}&\to\mathbb{R}^{s}\\ (x,t;\xi)&\mapsto(C_{L}\circ\sigma_{L-1}\circ\dots\circ\sigma_{1}\circ C_{1})(x,t;\xi)\end{aligned}\;\;:\;\;\theta\in\Theta\right\}.$
@@
(12) $C_{k}y=W_{k}y+b_{k},\quad\text{with}\quad W_{k}\in\mathbb{R}^{d_{k+1}\times d_{k}},\;\;b_{k}\in\mathbb{R}^{d_{k+1}},$
@@
$X_{\theta}(\xi)=g_{\theta}\left(\cdot;\xi\right),\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}.$
@@ 3.2 PCE-NN Based Neural Measure Spaces
To this end, consider a Polynomial Chaos Expansion (PCE)-type basis, [50], $\left\{\phi_{n}\right\}_{n\in\mathbb{N}}\subset L^{2}(\gamma)$ satisfying:
(13) $\forall q\in L^{2}(\gamma;\mathcal{H}),\quad\exists\,!\,\left\{a_{n}\right\}_{n\in\mathbb{N}}\subset\mathcal{H}:\quad q(x,t;\xi)=\sum_{n\in\mathbb{N}}a_{n}(x,t)\phi_{n}(\xi)\vskip-8.5359pt,$
@@
(14) $\mathcal{V}_{\mathcal{N}}\coloneqq\left\{g_{\theta}:\mathbb{R}^{d_{\text{in}}}\to\mathbb{R}^{d_{\text{out}}K}:\ g_{\theta}=g_{\theta}(x,t)\text{ for }\theta\in\Theta\right\},$
and the structure of the neural network functions $g_{\theta}(x,t)$ is as before, but now affecting only $x,t$ variables. Then, the corresponding form of neural measures is:
(15) $\mathcal{V}_{\mathcal{N}}^{*}=\left\{\left[X_{\theta}\right]_{\!\#}\gamma\in\mathcal{P}_{2}^{\gamma}\left(\,\,\mathcal{H}\,\,\right):\text{ where }X_{\theta}(\xi)=\sum_{n=1}^{K}g_{\theta}^{(n)}\,\phi_{n}(\xi)\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}\right\}.$
@@ 3.3 Galerkin-NN Based Neural Measure Spaces
Consider appropriate a given finite dimensional Galerkin subspace of $\mathcal{H}$, denoted by $\mathcal{H}_{h},$ let $M$ be its dimension and $\left\{\psi_{n}\right\}_{1}^{M}$ its basis. Then, we consider the elements of $L^{2}(\gamma;\mathcal{H})$
(16) $q(x,t;\xi)=\sum_{n=1}^{M}b_{n}(\xi)\psi_{n}(x,t).$
@@
$\mathcal{V}_{\mathcal{N}}\coloneqq\left\{g_{\theta}:\mathbb{R}^{d_{\text{ran}}}\to\mathbb{R}^{M}:g_{\theta}=g_{\theta}(\xi)\text{ for }\theta\in\Theta\right\},$
and the structure of the neural network functions $g_{\theta}(\xi)$ is as before, but now affecting only the $\xi$ variable. Then, the corresponding form of neural measures is:
@@ 4.1 Loss based on Wasserstein distance
Within the framework the present paper it is particularly convenient to consider Wasserstein distances. A key observation is that when one of the measures to be used in the loss is atomic, i.e., $\delta_{g},$ then the set of plans is reduced to $\mu\times\delta_{g},$ i.e., $\Pi(\mu,\delta_{g})=\{\mu\times\delta_{g}\}$, see (5.2.12) in [2]. Assuming $g\in\mathcal{H}$ and $\delta_{g}\in\mathcal{P}_{p}(\mathcal{H})$, we then have
(17) $\begin{split}W_{p}(\mu,\delta_{g})^{p}=&\inf_{\pi\in\Pi(\mu,\delta_{g})}\int_{\mathcal{H}}\int_{\mathcal{H}}\left\|x-y\right\|_{\mathcal{H}}^{p}\pi({\rm d}x,{\rm d}y)\\ =&\int_{\mathcal{H}}\int_{\mathcal{H}}\left\|x-y\right\|_{\mathcal{H}}^{p}\mu({\rm d}x)\delta_{g}({\rm d}y)\\ =&\int_{\mathcal{H}}\left\|x-g\right\|_{\mathcal{H}}^{p}\mu({\rm d}x)=\mathbb{E}_{X\sim\mu}\left\|X-g\right\|_{\mathcal{H}}^{p}.\end{split}$
We distinguish two cases. In case where $\mathcal{A}U(\xi)$ and $\mathcal{B}U(\xi)$ are constant with respect to $\xi$, we can easily check exists $f\in\tilde{\mathcal{H}}$ and $g\in\overline{\mathcal{H}}$ such that
@@
$\mu_{\theta^{*}}\in\operatorname*{arg\,min}_{\mu_{\theta}\in\mathcal{V}_{\mathcal{N}}^{*}}W_{p}(A\odot\mu_{\theta},\delta_{f})+W_{p}(B\odot\mu_{\theta},\delta_{g})$
@@
$\tilde{A}_{\xi}X_{\theta}(\xi)\coloneqq\mathcal{A}X_{\theta}(\xi)-\mathcal{A}U(\xi),\quad\text{ and }\quad\tilde{B}_{\xi}X_{\theta}(\xi)\coloneqq\mathcal{B}X_{\theta}(\xi)-\mathcal{B}U(\xi),$
@@
$\begin{split}d_{2,\gamma}\left(A\odot\nu_{u},\,A\odot\mu_{\theta}\right)^{2}+&d_{2,\gamma}\left(B\odot\nu_{u},\,B\odot\mu_{\theta}\right)^{2}\\ &=\left\|\mathcal{A}U-\mathcal{A}X_{\theta}\right\|_{L^{2}(\gamma;\tilde{\mathcal{H}})}^{2}+\left\|\mathcal{B}U-\mathcal{B}X_{\theta}\right\|_{L^{2}(\gamma;\overline{\mathcal{H}})}^{2}.\end{split}$
[…]
@@ 4.2 Other metrics and divergences
It is possible, however, to utilise alternative divergences or metrics to define the loss functional. We briefly review some potential choices below, see, eg., [7, 43, 33, 46, 6, 49, 44].
@@ 4.2.1 Divergences and Metrics: Definitions and Relationships
Divergences generalise metrics by dropping symmetry and the triangle inequality, while retaining non-negativity and vanishing only when the inputs are equal. A key family is the f-divergences: given probability measures $\mu\ll\nu$ and a convex function $f:\mathbb{R}_{+}\to\mathbb{R}\cup\left\{+\infty\right\}$ with $f(1)=0$, the $f$-divergence is defined as
@@
    (18) $d_{\mathcal{F}}(\mu,\nu)\coloneqq\sup_{f\in\mathcal{F}}\left|\mathbb{E}_{X\sim\mu}f(X)-\mathbb{E}_{Y\sim\nu}f(Y)\right|.$
@@
    Given a probability measure $\lambda\in\mathcal{P}(\mathcal{H})$, we define $\mathcal{P}_{\lambda}(\mathcal{H})\subset\mathcal{P}(\mathcal{H})$ to be the collection of probability measures that are absolutely continuous with respect to $\lambda$, i.e., $\mu\ll\lambda$. Let $\mu,\,\nu\in\mathcal{P}_{\lambda}(\mathcal{H})$ with Radon-Nikodym derivatives $f_{\mu},\,f_{\nu}$ with respect to $\lambda$. Then, TV distance between $\mu,\,\nu$ can be expressed as:
    $d_{TV}(\mu,\nu)\coloneqq\frac{1}{2}\left\|f_{\mu}-f_{\nu}\right\|_{L^{1}(\lambda)}.$
@@ 4.2.2 Approximating Distances in Finite-Dimensional Spaces
Given that our spaces, $V_{\Theta}=V_{\mathcal{N}}^{*}$, have a finite number of parameters, we can employ techniques developed in approximating distances of distributions in finite dimensions to compute the actual loss function. Indicative such methods include, GANs and computational optimal transport algorithms, [17, 36].
For approximating Wasserstein distances, one may use Sinkhorn Algorithm which introduces entropic regularisation for efficient computation and other approaches detailed in [36]. These approaches exploit the geometric and variational structure of optimal transport theory and are particularly relevant in generative modelling and distributional approximation tasks.
Generative Adversarial Networks (GANs) provide a flexible variational framework for approximating probability distributions by minimising statistical distances, notably Integral Probability Metrics (IPMs) and, under certain conditions, $f$-divergences.
@@
In the special case where the class of test functions $\mathcal{F}$ associated with a particular IPM (eg. as in equation Eq. 18) can be approximated by the class of discriminators $\overline{\mathcal{V}}_{\mathcal{N}}$. Then, this min-max formulation can be viewed as minimising IPM between $\mathbb{P}_{r}$ and generators distribution. Specifically, we define the class of neural measures:
[…]
@@
[…]
@@ 5 Numerical results
following [60]. This choice consistently yielded better convergence and expressiveness than the traditional hyperbolic tangent activation, especially for problems involving both periodic and non-periodic solutions. The parameter $a$ is specific to each layer and is optimized jointly with the network weights during training.
@@
All experiments were conducted using a unified codebase, which is publicly available at [3], ensuring reproducibility and ease of extension to other problem settings.
@@ 5.1 Bistable ordinary differential equation
(19) $\frac{\mathrm{d}u}{\mathrm{d}t}=-r(u-1)(2-u)(u-3),\quad t\in[0,8],$
@@
(20) $u(0)=u_{0}.$
The parameters $u_{0},r$ follow uniform distributions, $u_{0}\sim\mathcal{U}(0,4)$ and $r\sim\mathcal{U}(0.8,1.2)$. The equation has two stable ($u=1$, $u=3$) and one unstable ($u=2$) equilibrium solutions.
In Figure 4, we plot the training and testing loss functions throughout the optimisation. The spatial-temporal domain is resampled every 50 iterations, as are the parameters, which are drawn from their predefined uniform distribution. Each resampling uses 100 random samples from the domain. After each resampling step, the loss typically spikes due to the shift in training data but continues to decrease overall as training progresses. The testing loss begins to rise after around 400 iterations, indicating that the network starts to overfit the training data.
In Figure 1, we plot histograms of the solution to the bistable ODE Eq. 19. The reference solution is obtained using an implicit numerical solver. We observe good qualitative agreement between the reference and the PINN-based histograms. However, the Wasserstein distance between the two solutions increases over time. This is expected, as the solution’s distribution gradually concentrates into two Dirac delta peaks at $u=1$ and $u=3$. This sharp concentration makes convergence challenging, since the Wasserstein distance is highly sensitive to samples that have not fully collapsed onto the Dirac masses.
Figure 1: Histogram comparison at time $t=2$ of the solution of the bistable ODE Eq. 19 using an implicit numerical solver and the PINN architecture for different parameter values.